OpenStudy (anonymous):

Find the value of y in the mod 9 system. 5 x y = 4

OpenStudy (anonymous):

5y=4 y=4/5 what is mod 9 system if i may ask?

OpenStudy (cruffo):

\[5y=4 \;\text{mod(9)}\] ???

OpenStudy (anonymous):

modular system with a specific number of elements are analogous to the 12 oclock system

OpenStudy (cruffo):

There is an algebraic way to solve these, but since there are only 9 elements, trial and error can also work. 5(1) = 5 mod(9) 5(2) = 10 = 1 mod(9) 5(3) = 15 = 6 mod(9) 5(4) = 20 = 2 mod(9) 5(5) = 25 = 7 mod(9) 5(6) = 30 = 3 mod(9) 5(7) = 35 = 8 mod(9) 5(8) = 40 = 4 (mod(9) ------ solution y = 8

OpenStudy (anonymous):

made easier to understand you writing out...thx

OpenStudy (anonymous):

what is that?

OpenStudy (anonymous):

Modulo congruence \[a \cong b (mod\ c) \implies \exists k \in Z\ |\ {a-b \over c} = k \]

OpenStudy (cruffo):

multiply, then divide the result by the indicated mod #, take the remainder. Most likely studied in beginning number theory, or abstract algebra, or cryptography.

OpenStudy (anonymous):

Are you looking for solutions where 5y is congruent to 4 ? Or are you looking for solutions where 5y (mod 9) = 4?

OpenStudy (anonymous):

Actually those would be the same since 4 mod 9 is 4.

OpenStudy (cruffo):

:) I was just gonna ask about that!